General Meeting Summary 6/9/22

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Paulo, Tanja, Vardan, Vladimir, Aaron, Hamlet, Brad, Rolf, Bogdan, Abishek, Pramita, Carlos, Alexandre, Julie, Hakob, Marie, David H


CRYTAL MODULE PREPARATION (Vladimir)

  • Team of 7 people (Vladimir, Abishek, Hamlet, Josh, Petr, Vardan, Zheng) from 4 institutions (AANL, CUA, UIC, MSU)
  • Completed the crystal module preparation during the month of May
  • 1100 crystals
  • team worked every day of the month for the entire day
  • All but 100 crystals were prepared in RM 126 - 100 crystals prepared in NPS cleanroom
  • Crystals were stored on shelves in the NPS assembly area
  • all labeled from 1-1100


CRYSTAL STACKING (Carlos)

  • Frame prepared
  • 19 rows out of 36 were stacked, expect other half to be ready by Friday afternoon
  • Next step: PMT assembly
  • done column by column - all cables installed on each PMT
  • Reason for not fully parallelizing PMT assembly with stacking: one crate that was sent from Orsay was delayed and only received this week - low voltage and some tools needed for assembly were in it
  • Next week: start cabling - anticipate cabling complete for a few columns before Orsay group leaves
  • Orsay team can return in September to finish cabling or for further tests (cooling, electronics, etc.)
  • Action items:
  • organize meeting on Thursday 6/16 at 9AM ET to discuss path forward, e.g., cabling after Orsay team leaves, possible detector checkout, etc.


MECHANICAL DESIGN/ENGINEERING (Paulo)

  • SHMS/NPS installed/fitup
  • platform completely installed and removed
  • new detector deck section installed permanently
  • Detector frame assembled
  • Detector box installed onto detector frame and detector assembly ongoing by Orsay team
  • lifting fixture expected in July
  • beamline 1 (run group 1 - go to small angle) under physcis review and ready for fabrication
  • beamline 2 (run group 2 - reduce radiation) under physics review, some changes requested, and ready for fabrication
  • Items left to do:
  • target pivot reinforcement
  • cabling and cable management
  • Action items:
  • check on if any magnetic items in NPS frame


UPDATES FROM DSG (Aaron)

  • CAEN HV Controls and Monitoring
  • Labview hardware interlock program development
  • temperature and humidity sensor
  • dew point
  • etc.
  • LED graphical programming ongoing
  • HV cables
  • tested and podded
  • 21 ready for cabling - can be moved from 126 to 108
  • Action items:
  • check on HV cables - ok to move cables from 126 to 108
  • LED graphical programming


DAQ/ELECTRONICS AND SOFTWARE (David H, Brad)

  • Main effort centered on updating HCANA to be compatible with MOD10 (full waveform) data
  • data from PionLT used for setting up and testing
  • Other efforts
  • replay script
  • database
  • Moving DAQ/electronics equipment into RM108 for setting up the DAQ
  • HV and power supplies
  • HV cables are in 126
  • power supplies etc in ESB
  • LV power supply in 126 along with carts
  • cabling and connectors on site - some cables still need to be assembled
  • LEDs on each PMTs - LED boards are ready
  • Need good oscilloscope in 108 - there is one available and planned to be moved there
  • Action items:
  • continue checking off items from list


GENERAL DISCUSSION

  • Action item: make checklist to review action items to make sure that nothing forgotten
  • impact on NPS magnet on HMS optics
  • NPS frame magnetic/nonmagnetic
  • looks like aluminum
  • check if steel (bolts etc.)
  • seems all on frame is nonmagnetic
  • Mailing lists have been setup: nps and nps-daq mailing lists
  • no action needed for current collaborators
  • if new students/collaborators please sign up through JLab mailing lists
  • Talk at Hall A/C meeting on June 16th and June 17th
  • Propose one 40min long talk on both detector and physics
  • start the talk with assembly acivities to show all the progress
  • talk might have ~10 slides on physics upcoming and new proposed (ask experiments to provide one slide each - there will also be full talks on these proposals later in the workshop)
  • speaker: Vladimir Berdnikov
  • Talk with focus on new opportunities with NPS (DDVCS, etc.) at Workshop at VTech on July 18-22
  • Abstract submission still open


NEXT MEETING: THURSDAY JUNE 30TH AT 9AM ET